How can I differentiate 99x with respect to x pls get me the answer as soon as possible dear sir/madam

Grade:11

1 Answers

Mahima Kanawat
6 years ago
Dear Student .....The formula for differentiating nx is nxⁿ-¹....here , n=99 Hence ..d(99x)/dx=99x¹-¹=99x°=99×1=99...With Regards ...Mahima...(Askiitians Forum Expert )